OpenStudy (anonymous):

Simplify. A. –32 B. –32xy C. 32xy D. 32

OpenStudy (anonymous):

simplify wat

OpenStudy (anonymous):

-3/4{-8x}4/3{4y}{1}

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

\(-3/4(-8x)4/3(4y)(1)\) Multiply: \(-3/4(-8x)4/3(4y)(1)\) \((6x)(4/3)(4y)(1)\) \((8x)(4y)(1)\) \(32xy \times 1\) Can you multiply that? @newwar248

OpenStudy (anonymous):

@Jackemup98

OpenStudy (igreen):

Can you multiply 32xy * 1? That'll be your answer.

OpenStudy (anonymous):

@kath_010

OpenStudy (igreen):

Anything multiplied 1 is the same thing it was before. Examples: 123 * 1 = 123 814325 * 1 = 814325 9 * 1 = 9 235x * 1 = 235x

OpenStudy (igreen):

So what's 32xy * 1?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@SnakeEyes_417

OpenStudy (igreen):

32xy * 1 = 32xy, so the answer is 32xy. @newwar248

OpenStudy (anonymous):

@Shivam1115

OpenStudy (anonymous):

thanks